((D)^2 - (D')^2 - 3D + 3D')z = sin(x-2y)
1
Expert's answer
2021-09-16T06:54:52-0400

((D)2-(D')2-3D+3D')z=Sin(x-2y)

Auxiliary equation of(D2-D'2-3D+3D')z=0 is;

m2-12-3m+3*1=0

m2-3m+2=0

m2-2m-m+2=0

m(m-2)-1(m-2)=0

(m-2)(m-1)=0

m=2 or 1

Complementary function is "\phi_1(y+x)+\phi_2(y+2x)"

P.I="\frac{1}{D^{2}-D'^{2}-3D+3D'}Sin(x-2y)"

Let f(D,D')=D2-D'2-3D+3D'

Now co-eff of x is 1=a(say)

And that of y is -2=b(say)

so that f(a,b)=f(1,-2)

=12-(-2)2-3(1)+3(-2)

=1-4-3-6

=-12 "\not=" 0

P.I="\frac{1}{f(a,b)}\iint Sin \ vdvdv" let x-2y=v

="\frac{-1}{12}\int -Cot \ vdv=\frac{1}{12}Sin\ v"

="\frac{1}{12}Sin(x-2y)"

Therefore the general solution is :

C.F+P.I

"=\phi_1(y+x)+\phi_2(y+2x)+\frac{1}{12}Sin(x-2y)"